OpenStudy (anonymous):

Solve 4x^2=28 by finding square roots. Please show me how to solve this by including the steps.

OpenStudy (zzr0ck3r):

first divide both sides by 4, and tell me what you have.

OpenStudy (anonymous):

x^2=7

OpenStudy (zzr0ck3r):

ok now take the square root of both sides and tell me what you have

OpenStudy (anonymous):

x= 2.64

OpenStudy (anonymous):

?

OpenStudy (zzr0ck3r):

well, you should have gotten \(x=\pm \sqrt{7}\) So \(x=\sqrt{7}\) or \(-\sqrt{7}\)

OpenStudy (anonymous):

Thanks :)